Udit received $900 from his parents as a graduation present. He invested part of it at 2% interest and the remainder at 10%. If the total yearly interest amounted to $50, how much did he invest at each rate?
Answer by TimothyLamb(4379)

x = at 2%
y = at 10%
---
x + y = 900
0.02x + 0.10y = 50
---
put the system of linear equations into standard form
---
x + y = 900
0.02x + 0.10y = 50
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = at 2% = 500
y = at 10% = 400
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---